OpenStudy (anonymous):

Indefinite integration: 2x (x^2 + 1) ^4 dx

OpenStudy (anonymous):

u = x^+1

OpenStudy (anonymous):

du = 2x dx

OpenStudy (anonymous):

dx= du/2x So, u^4 du = u^5/5

OpenStudy (anonymous):

or, 1/5(x^2+1)^5

OpenStudy (anonymous):

Thanks!